OpenStudy (anonymous):

What is the second derivative of f(x) = x/ (x^2-1)

OpenStudy (anonymous):

I got the first derivative as -x^2-1 / (x^2-1)^2 but i dont know how to do the 2nd derivative

myininaya (myininaya):

Did you try a sub-y :)

OpenStudy (anonymous):

I need to do it the quotient rule way

myininaya (myininaya):

sorry mis-read

myininaya (myininaya):

So Just apply quotient rule again

OpenStudy (anonymous):

i got it till -2x(x^2-1)^2 - 4x^3 - 4x / (x^2-1)^4 it looks wrong to me..so i dont know

myininaya (myininaya):

\[f'=\frac{-x^2-1}{(x^2-1)^2}\] \[f''=\frac{(-2x)(x^2-1)^2-(-x^2-1)(2 \cdot (x^2-1) \cdot (2x)}{(x^2-1)^4} \]

myininaya (myininaya):

Did you get this far correct?

sam (.sam.):

Final answer should be \[\frac{d^2}{dx^2}\left(\frac{x}{x^2-1}\right)=\frac{2 x \left(x^2+3\right)}{\left(x^2-1\right)^3}\]

OpenStudy (anonymous):

why is there 2 (-x^2 -1)? there should be only 1 of them on the numerator isnt it

OpenStudy (anonymous):

what...i got like -2x(x^4 +3) / (x^2 -1 )^4...

myininaya (myininaya):

Hey cmkc I left the top the same and used chain rule for the bottom

OpenStudy (anonymous):

ohhh i know what i did wrong!! thank you both of u!!

myininaya (myininaya):

For that part of the quotient rule anyways

OpenStudy (anonymous):

yea, forgot to put the initial ..haha thanks so much!!

OpenStudy (anonymous):

I still got it wrong can u check?

sam (.sam.):

wait

OpenStudy (anonymous):

Ok thanks! I will be waiting :)

sam (.sam.):

\[y=\frac{x}{x^2-1}\] \[y'=\frac{(x^2-1)-x(2x)}{(x^2-1)^2}\] \[y'=\frac{-x^2-1}{(x^2-1)^2}\] \[y''=\frac{(x^2-1)^2(-2x)-(-x^2-1)(2)(x^2-1)(2x)}{(x^2-1)^4}\] \[y''=\frac{-2x(x^2-1)^2-(-2x^3-2x)(2)(x^2-1)}{(x^2-1)^4}\] \[y''=\frac{(x^2-1)[2x(x^2-1)+4x^3+4x]}{(x^2-1)^4}\] \[y''=\frac{[-2x^3+2x+4x^3+4x]}{(x^2-1)^3}\] \[y''=\frac{2x^3+6x}{(x^2-1)^3}\] \[y''=\frac{2x(x^2+3)}{(x^2-1)^3}\]

OpenStudy (anonymous):

thank u thank u

sam (.sam.):

no prob